OpenStudy (anonymous):

linear equations how to do -3x+y=2?

OpenStudy (nincompoop):

y-intercept and x-intercept

OpenStudy (anonymous):

how do i solve for them?

OpenStudy (anonymous):

First, solve for y. The y intercept will be where x=o. Therefore, plug 0 into your new equation and this should yield the y intercept. Likewise, if you plug zero in for y, you should get the x intercept.

OpenStudy (nincompoop):

y = mx+b -3x + y = 2 (rewrite it to y-intercept) y= 3x+2 solve for y-intercept by changing the value of x to zero y= 3(0) + 2 y = 2 solve for x-intercept by changing the value of y to zero do the rest :)

OpenStudy (anonymous):

so for x is -3(0)+y=2?

OpenStudy (anonymous):

or for y do i leave it as 2 in it then solve for x?

OpenStudy (nincompoop):

it will be: y = mx+b y= 3x + 2 (solve for x-intercept by changing the value of y to zero) 0 = 3x + 2 -3x = 2 (solve for x by dividing the coefficient with -3) x = -2/3

OpenStudy (nincompoop):

values: x = -2/3 y = 2

OpenStudy (anonymous):

if I was to do this for a different problem would I do the same thing?

OpenStudy (anonymous):

yes! Just make sure to solve for y first so you have y=mx+b

OpenStudy (nincompoop):

yes :) when you are given relatively same equation, just always solve the y-intercept and x-intercept. note that y-intercept means the part in the cartesian plane where the line touches the y-axis which, the value of x is zero. And for the x-intercept, it would be the part of the cartesian plane where the line touches the x-axis, which the value of y is zero.

OpenStudy (anonymous):

can you give me a example and I try to do it and see if I get it?

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

how would i graph -2/3 on a graph? it doesnt have that low

OpenStudy (anonymous):

it doesn't support it

OpenStudy (anonymous):

oh i see never mind thank you.